For the function f(x) 52-7, evaluate and fully simplify each of the following. Preview f(a + h) - f() Preview
media%2F3f8%2F3f865e1b-4ace-466b-a921-2a
Answer #1

1)
f(x) = 5x^2 - x - 7

a)
f(x+h) = 5(x+h)^2 - (x+h) - 7
= 5(x^2 + h^2 + 2hx) - x - h - 7
= 5x^2 + 5h^2 + 10hx - x - h - 7

b)
(f(x+h) - f(x))/h
= ((5x^2 + 5h^2 + 10hx - x - h - 7) - (5x^2 - x - 7))/h
= (5h^2 + 10hx - h ) / h
= 5h + 10x - 1

2)
g(x) = -6x + 4

a)
g(-9) = -6*(-9) + 4
= 54 + 4
= 58

b)
g(x) = 52
-6x + 4 = 52
6x = -48
x = -8